OpenStudy (anonymous):

Lol i still can't get this halp 4(x + 1) = 4x + 4. 5 = 4 x 2 + 4 5 = 12

OpenStudy (anonymous):

okay so i assume you just want to solve for x in the function 4(x + 1) = 4x + 4. right?

OpenStudy (anonymous):

oops the 2 second lines are my work that i messed up on the problem is 4(x+1) = 4x + 4

OpenStudy (anonymous):

Okay do you remember how to distribute the 4 into (x+1)?

OpenStudy (anonymous):

the options for it are no solution 0 all real numbers 1

OpenStudy (anonymous):

It's like taking 4 and multiplying x and +1 separately. Just think of it like that when you get started.

OpenStudy (anonymous):

Did you get that @misty_fox?

OpenStudy (anonymous):

4(x+1) = 4x + 4 we start of by distributing 4(x+1) = 4(x + 1) and here you can see that everything cancels out which usually means the solution is all real numbers to make sure you can plug in both 0 and 1 just to make sure and sure enough they both work

OpenStudy (anonymous):

woops ment to say factoring not distributing but you can distribute too and you'd still get to equal functions on either side

OpenStudy (anonymous):

Yes watchpencilpaper is right! :)

OpenStudy (mathstudent55):